# Transverse and longitudinal spin alignment from color fields in heavy ion collisions

### v1

@@ Abstract
We analyze the spin alignment stemming from spin correlation of the quark and antiquark induced by background color fields in relativistic heavy ion collisions. The quark-coalescence equation relating the collision kernel of the vector-meson kinetic equation to spin alignment is expanded to the relativistic case. Focusing on the color-octet contribution, the spin alignment for $\phi$ mesons from glasma fields with momentum dependence is investigated, where the scenarios for different spin quantization axes are considered. Moreover, we qualitatively analyze the spin alignment from isotropic color fields in the quark gluon plasma phase for comparison. In particular, we propose that the experimental measurement of spin alignment along the beam direction, dubbed as the longitudinal spin alignment, could be useful to identify the dominance of longitudinal spin correlation potentially led by the glasma effect in high-energy nuclear collisions.
@@ I Introduction
The paper is structured as follows: In Sec. II, we briefly review the derivation of axial Wigner functions dictating the spin polarization of quarks and antiquarks from the quantum kinetic theory (QKT) with background color fields. In Sec. III, we revisit the quark coalescence scenario delineated by the collision kernel of the kinetic theory of vector mesons and derive a more general expression, especially suitable for studying the glasma effect, beyond the non-relativistic approximation and equal-mass condition. The choices of different spin quantization axes are also discussed. In Sec. IV, we evaluate the spin alignment of $\phi$ mesons from the glasma effect and show the numerical results with momentum dependence. In Sec. V, we further analyze the spin alignment of $\phi$ mesons from color fields in QGP with an emphasis on qualitative features and the comparison with the results from glasma fields. Finally, we present conclusions and outlook in Sec VI. Some technical details have been relegated to the appendices.
@@ II Spin transport of quarks with color fields
In this section, we briefly review the QKT for spin transport of quarks with color fields constructed in Refs. [51, 52] and the approximate solution for dynamical spin polarization obtained in Ref. [60, 59]. For spin transport, one may focus on the vector and axial-vector components of Wigner functions for massive fermions, $\mathcal{V}^{\mu}(p,x)$ and $\mathcal{A}^{\mu}(p,x)$, where the former and the latter are responsible for the particle-number and spin-polarization spectra, respectively. The dynamical evolution of $\mathcal{V}^{\mu}(p,x)$ and $\mathcal{A}^{\mu}(p,x)$ in phase space can then be tracked by utilizing the QKT for relativistic fermions [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86] (see also a recent review  [87] and references therein) with further inclusion of the color degrees of freedom [51, 52, 88]. For quarks carrying color charges, they can be further decomposed into the color-singlet and color-octet components,
@@
$\mathcal{A}^{{\rm s}\mu}(\bm{p},x)$$\equiv$ $\int\frac{dp_{0}}{2\pi}\mathcal{A}^{{\rm s}\mu}(p,x)=\frac{1}{2\epsilon_{\bm{p}}}\Big{[}\tilde{a}^{{\rm s}\mu}-\frac{\hbar}{4N_{c}}\tilde{F}^{a\mu\nu}\Big{(}\partial_{p\nu}f^{a}_{\rm V}-\frac{\epsilon_{\bm{p}}}{2}\partial_{p_{\perp}\nu}(f^{a}_{V}/\epsilon_{\bm{p}})\Big{)}\Big{]}_{p_{0}=\epsilon_{\bm{p}}},$
@@
$=$ $\frac{1}{2\epsilon_{\bm{p}}}\Big{[}\tilde{a}^{a\mu}-\frac{\hbar}{2}\tilde{F}^{a\mu\nu}\Big{(}\partial_{p\nu}f^{\rm s}_{V}-\frac{\epsilon_{\bm{p}}}{2}\partial_{p_{\perp}\nu}(f^{\rm s}_{V}/\epsilon_{\bm{p}})\Big{)}$
$-\frac{\hbar d^{bca}}{4}\tilde{F}^{b\mu\nu}\Big{(}\partial_{p\nu}f^{c}_{V}-\frac{\epsilon_{\bm{p}}}{2}\partial_{p_{\perp}\nu}(f^{c}_{V}/\epsilon_{\bm{p}})\Big{)}\Big{]}_{p_{0}=\epsilon_{\bm{p}}},$
@@
The dynamics of vector-charge distribution functions and effective spin four vector in phase space are governed by the coupled scalar kinetic equations (SKE),
@@
$p^{\rho}\partial_{\rho}\tilde{a}^{\rm s\mu}+\frac{g}{2N_{c}}\Big{(}p^{\rho}F^{a}_{\nu\rho}\partial_{p}^{\nu}\tilde{a}^{a\mu}+F^{a\nu\mu}\tilde{a}^{a}_{\nu}\Big{)}-\frac{\hbar}{4N_{c}}\epsilon^{\mu\nu\rho\sigma}p_{\rho}(\partial_{\sigma}gF^{a}_{\beta\nu})\partial_{p}^{\beta}f_{V}^{a}=\mathcal{C}^{\mu}_{\rm s},$ (7)
@@
where $D_{\rho}O^{a}\equiv\partial_{\rho}O^{a}-gf^{bca}A^{b}_{\rho}O^{c}$ denotes the covariant derivative for an arbitrary color object $O^{a}$. Here we do not specify the collision terms explicitly. See Refs. [51, 52] for more detailed explanations of the theoretical framework.
@@
For the glasma effect on spin alignment proposed in Ref. [60], one may further drop the contributions from late-time fields by assuming $B^{ai}(t_{\rm f}),\,E^{ai}(t_{\rm f})\ll B^{ai}(t_{\rm i}),\,E^{ai}(t_{\rm i})$ for time-decreasing glasma fields. We shall regard $\tilde{a}^{a\mu}(p,x)|_{t_{\rm f}=t_{\rm th}}$ as the early-time polarization induced in the glamsa phase ending at $t_{\rm th}$ as the thermailzation time and the onset of the QGP phase. For $t>t_{\rm th}$, the glasma fields vanish and the collisional effect in QGP should result in the spin relaxation [^2] The sharp transition between the glasma phase and QGP phase is assumed for simplicity, which may not reflect the realistic situation with an intermediate phase for thermalization.. Then one has to solve
@@
Neglecting the non-dynamical contribution corresponding to the part besides $\tilde{a}^{a\mu}$ in $\mathcal{A}^{ai}$ due to the absence of color fields in late times, we arrives at
@@ III.1 Generalization of the coalescence scenario
$\quad+2{\rm Re}\big{(}k\cdot\epsilon(\lambda,{\bm{q}})\mathcal{A}_{\bar{q}}(\bm{p}^{\prime},x)\cdot\epsilon^{*}(\lambda,{\bm{q}})\big{)}k\cdot\mathcal{A}_{q}(\bm{p},x)\Big{)}\Bigg{\}}\Bigg{|}_{p,p^{\prime}}$ (21)
@@
We also use subscripts $q$ and $\bar{q}$ to represent the contributions from quarks nd antiquarks for the vector distribution functions and axial-vector Wigner functions. We then work in the vector-meson rest frame, $q^{\mu}=(M,{\bm{0}})$, such that $\epsilon^{\mu}(\lambda,0)=(0,\bm{\epsilon}_{\lambda})\equiv\epsilon_{\lambda}^{\mu}$. Accordingly, the onshell and kinematic conditions give rise to
@@
Based on the complicated yet straightforward computations shown in Appendix. A, it is found
@@
Here we have dropped $f^{a}_{Vq,\bar{q}}$-related terms as the higher-order corrections for the polarization independent part and simply denoted $f_{Vq,\bar{q}}=f^{\rm s}_{Vq,\bar{q}}$ for brevity. Despite the assumption for a particular form of the axial Wigner functions involved, Eq. (29) now serves as a general expression satisfying the momentum conservation and applicable for $m_{q}\neq m_{\bar{q}}$.
@@ IV.3 Numerical results
In Fig. (2), we plot $\rho^{\rm o}_{00}$ with $y_{q}$ dependence at fixed $\phi$ and $q_{T}$. Generically, the $y_{q}$ dependence is relatively mild. Basically, $\rho^{\rm o}_{00}$ monotonically increases for $q_{x}=q_{y}$ and decreases for $q_{y}=0$. For large $q_{T}$, $\rho^{\rm o}_{00}$ could be larger than $1/3$ for $q_{y}=0$ as already indicated Fig. 1. We further show contour plots for $\delta\rho^{\rm o}_{00}$ with respect to transverse momenta in Fig. (3). In general, $\rho^{\rm o}_{00}<1/3$ in most of kinematic regions.
@@
Analogously, we may also numerically evaluate $\rho^{\rm i}_{00}$ for the in-plane spin alignment for $\phi$ mesons. Due to symmetry, the results for $\rho^{\rm i}_{00}$ with $q_{T}$ dependence at fixed $\phi$ and $y_{q}$ are the same as those shown in Fig. (1) by exchanging the results for $\phi=0$ and $\phi=\pi/2$. In Fig. 4, we further present $\delta\rho^{\rm i}_{00}$ with $\phi$ dependence and $\rho^{\rm i}_{00}$ with $y_{q}$ dependence, respectively. In particular, for the large-momentum limit, we simply need to exchange $\hat{v}_{x}$ and $\hat{v}_{y}$ in Eq. (86), which yields $\delta\rho^{\rm o}_{00}\sim-(2-3\sin^{2}\phi)$, and the qualitatively features of $\delta\rho^{\rm i}_{00}$ with $\phi$ dependence at large $q_{T}$ are understood. Also, $\rho^{\rm i}_{00}<1/3$ for either $\phi=\pi$ or $\phi=0$ is expected. The $y_{q}$ dependence is generically mild. The contour plots in Fig. 5 correspond to those in Fig. 3 with $90$-degree rotations.
@@ V Spin alignment from color fields in QGP
$\rho_{00}(q)=\frac{1-\big{[}\big{(}\tilde{C}_{B1}+\tilde{C}_{B2}\big{)}\chi^{yy}_{B}+\tilde{C}_{B1}(\chi^{xx}_{B}+\chi^{zz}_{B}\big{)}+\big{(}\tilde{C}_{E1}+\tilde{C}_{E2}\big{)}\chi^{yy}_{E}+\tilde{C}_{E1}(\chi^{xx}_{E}+\chi^{zz}_{E}\big{)}\big{]}}{3-\big{[}\big{(}3\tilde{C}_{B1}+\tilde{C}_{B2}\big{)}\chi_{B}+\big{(}3\tilde{C}_{E1}+\tilde{C}_{E2}\big{)}\chi_{E}\big{]}},$ (87)
@@ Acknowledgements.
The author would like to thank the hospitality of ECT* during the workshop “Spin and Quantum Features of QCD Plasma” and other participants, especially A. Tang, for helpful discussions. This work is supported by National Science and Technology Council (Taiwan) under Grants No. MOST 110-2112-M-001-070-MY3 and No. NSTC 113-2628-M-001-009-MY4.

### v2

@@ Abstract
We analyze the spin alignment of vector mesons stemming from spin correlation of the quark and antiquark induced by background color fields in relativistic heavy ion collisions. The quark-coalescence equation relating the collision kernel of the vector-meson kinetic equation to spin alignment is expanded to the relativistic case. Focusing on the color-octet contribution, the spin alignment for $\phi$ mesons from glasma fields with momentum dependence is investigated, where the scenarios for different spin quantization axes are considered. Moreover, we qualitatively analyze the spin alignment from isotropic color fields in the quark gluon plasma phase for comparison. In particular, we propose that the experimental measurement of spin alignment along the beam direction, dubbed as the longitudinal spin alignment, could be useful to identify the dominance of longitudinal spin correlation potentially led by the glasma effect in high-energy nuclear collisions.
@@ I Introduction
The paper is structured as follows: In Sec. II, we briefly review the derivation of axial Wigner functions dictating the spin polarization of quarks and antiquarks from the quantum kinetic theory (QKT) with background color fields. In Sec. III, we revisit the quark coalescence scenario delineated by the collision kernel of the kinetic theory of vector mesons and derive a more general expression, especially suitable for studying the glasma effect, beyond the non-relativistic approximation and equal-mass condition. The choices of different spin quantization axes are also discussed. In Sec. IV, we evaluate the spin alignment of $\phi$ mesons from the glasma effect and show the numerical results with momentum dependence. In Sec. V, we further analyze the spin alignment of $\phi$ mesons from color fields in QGP with an emphasis on qualitative features and the comparison with the results from glasma fields. Finally, we present conclusions and outlook in Sec VI. Some technical details have been relegated to the Appendix.
@@ II Spin transport of quarks with color fields
In this section, we briefly review the QKT for spin transport of quarks with color fields constructed in Refs. [51, 52] and the approximate solution for dynamical spin polarization obtained in Refs. [60, 59]. For spin transport, one may focus on the vector and axial-vector components of Wigner functions for massive fermions, $\mathcal{V}^{\mu}(p,x)$ and $\mathcal{A}^{\mu}(p,x)$, where the former and the latter are responsible for the particle-number and spin-polarization spectra, respectively. The dynamical evolution of $\mathcal{V}^{\mu}(p,x)$ and $\mathcal{A}^{\mu}(p,x)$ in phase space can then be tracked by utilizing the QKT for relativistic fermions [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86] (see also a recent review  [87] and references therein) with further inclusion of the color degrees of freedom [51, 52, 88]. For quarks carrying color charges, they can be further decomposed into the color-singlet and color-octet components,
@@
$\mathcal{A}^{{\rm s}\mu}(\bm{p},x)$$\equiv$ $\int\frac{dp_{0}}{2\pi}\mathcal{A}^{{\rm s}\mu}(p,x)=\frac{1}{2\epsilon_{\bm{p}}}\Big{[}\tilde{a}^{{\rm s}\mu}-\frac{\hbar g}{4N_{c}}\tilde{F}^{a\mu\nu}\Big{(}\partial_{p\nu}f^{a}_{\rm V}-\frac{\epsilon_{\bm{p}}}{2}\partial_{p_{\perp}\nu}(f^{a}_{V}/\epsilon_{\bm{p}})\Big{)}\Big{]}_{p_{0}=\epsilon_{\bm{p}}},$
@@
$=$ $\frac{1}{2\epsilon_{\bm{p}}}\Big{[}\tilde{a}^{a\mu}-\frac{\hbar g}{2}\tilde{F}^{a\mu\nu}\Big{(}\partial_{p\nu}f^{\rm s}_{V}-\frac{\epsilon_{\bm{p}}}{2}\partial_{p_{\perp}\nu}(f^{\rm s}_{V}/\epsilon_{\bm{p}})\Big{)}$
$-\frac{\hbar gd^{bca}}{4}\tilde{F}^{b\mu\nu}\Big{(}\partial_{p\nu}f^{c}_{V}-\frac{\epsilon_{\bm{p}}}{2}\partial_{p_{\perp}\nu}(f^{c}_{V}/\epsilon_{\bm{p}})\Big{)}\Big{]}_{p_{0}=\epsilon_{\bm{p}}},$
@@
As constructed from the Wigner-function approach in Refs. [51, 52], the dynamics of vector-charge distribution functions and effective spin four vector in phase space are governed by the coupled scalar kinetic equations (SKE),
@@
$p^{\rho}\partial_{\rho}\tilde{a}^{\rm s\mu}+\frac{g}{2N_{c}}\Big{(}p^{\rho}F^{a}_{\nu\rho}\partial_{p}^{\nu}\tilde{a}^{a\mu}+F^{a\nu\mu}\tilde{a}^{a}_{\nu}\Big{)}-\frac{\hbar}{4N_{c}}\epsilon^{\mu\nu\rho\sigma}p_{\rho}(D_{\sigma}gF^{a}_{\beta\nu})\partial_{p}^{\beta}f_{V}^{a}=\mathcal{C}^{\mu}_{\rm s},$ (7)
@@
where $D_{\rho}O^{a}\equiv\partial_{\rho}O^{a}-gf^{bca}A^{b}_{\rho}O^{c}$ denotes the covariant derivative for an arbitrary color object $O^{a}$. Here we do not specify the collision terms, $\mathcal{C}_{\rm s}$, $\mathcal{C}^{a}_{\rm o}$, $\mathcal{C}^{\mu}_{\rm s}$, and $\mathcal{C}^{a\mu}_{\rm o}$, explicitly, which depend on the details of interaction for the considered systems. For phenomenological purpose, we will later adopt the relaxation-time approximation for these collision terms in the QGP phase as the previous study in Ref. [60]. See Refs. [51, 52] for more detailed explanations of the theoretical framework and Ref. [60] for phenomenological applications.
@@
For the glasma effect on spin alignment proposed in Ref. [60], one may further drop the contributions from late-time fields by assuming $B^{ai}(t_{\rm f}),\,E^{ai}(t_{\rm f})\ll B^{ai}(t_{\rm i}),\,E^{ai}(t_{\rm i})$ for time-decreasing glasma fields. We shall regard $\tilde{a}^{a\mu}(p,x)|_{t_{\rm f}=t_{\rm th}}$ as the early-time polarization induced in the glamsa phase ending at $t_{\rm th}$ as the thermalization time and the onset of the QGP phase. For $t>t_{\rm th}$, the glasma fields vanish and the collisional effect in QGP should result in the spin relaxation [^2] The sharp transition between the glasma phase and QGP phase is assumed for simplicity, which may not reflect the realistic situation with an intermediate phase for thermalization.. Then one has to solve
@@
Neglecting the non-dynamical contribution corresponding to the part besides $\tilde{a}^{a\mu}$ in $\mathcal{A}^{ai}$ due to the absence of color fields in late times, we arrive at
@@
In the following sections, we will evaluate the spin alignment from the spin correlation of quarks and antiquarks induced by the correlated background color fields in the quark coalescence scenario. Due to technical complications, here we briefly summarize the general strategy and setup. Since the spin alignment of vector mesons is measured in the rest frame of vector mesons, we solve the polarization-dependent kinetic theory for vector mesons in the rest frame with the collision term of quark coalescence shown in Ref. [60]. The polarization dependence of such a collision term originates from the correlations of background color fields encoded in the axial Wigner functions of quarks and antiquarks obtained above. However, the correlations of color fields in the rest frame implicitly carry momentum dependence of the vector mesons in the lab frame. Such a momentum dependence is essential to yield anisotropic spin correlation for spin alignment when background color fields are isotropic in the lab frame. Moreover, the correlation functions of color fields are usually computed in the lab frame. We hence further make Lorentz transformation of the color-field correlators in the vector-meson rest frame to the lab frame in order to retrieve the momentum dependence of the spin-alignment spectrum for vector mesons. For a practical application to the spin alignment of $\phi$ mesons, we consider both the dynamical spin alignment led by anisotropic color fields from the glasma phase as an initial-state effect and the non-dynamical one triggered by isotropic color fields characterizing the interaction with soft gluons in the QGP phase as a final-state effect. In the former case, the key assumptions are first the strange quarks and antiquarks are produced in early times with strong glasma fields and second the spin relaxation time is relatively slow as opposed to the rapid thermalization of quarks and antiquarks that reach thermal equilibrium at the outset of QGP phase. For the latter case, we only consider the impact of constant background color fields on the spin correlation just before the quark coalescence for simplicity and conduct a qualitative analysis for spin alignment.
@@ III.1 Generalization of the coalescence scenario
$\quad+2{\rm Re}\big{(}k\cdot\epsilon(\lambda,{\bm{q}})\mathcal{A}_{\bar{q}}(\bm{p}^{\prime},x)\cdot\epsilon^{*}(\lambda,{\bm{q}})\big{)}k\cdot\mathcal{A}_{q}(\bm{p},x)\Big{)}\Bigg{]}\Bigg{\}}\Bigg{|}_{p,p^{\prime}}$ (21)
@@
We also use subscripts $q$ and $\bar{q}$ to represent the contributions from quarks and antiquarks for the vector distribution functions and axial-vector Wigner functions. We then work in the vector-meson rest frame, $q^{\mu}=(M,{\bm{0}})$, such that $\epsilon^{\mu}(\lambda,0)=(0,\bm{\epsilon}_{\lambda})\equiv\epsilon_{\lambda}^{\mu}$. Accordingly, the onshell and kinematic conditions give rise to
@@
Based on the complicated yet straightforward computations shown in the Appendix, it is found
@@
The notation $\langle O^{a}\,O^{b}\rangle$ represents a quantum expectation value of the correlation function for field operators $O^{a}$ and $O^{b}$. Here we have dropped $f^{a}_{Vq,\bar{q}}$-related terms as the higher-order corrections for the polarization independent part and simply denoted $f_{Vq,\bar{q}}=f^{\rm s}_{Vq,\bar{q}}$ for brevity. Despite the assumption for a particular form of the axial Wigner functions involved, Eq. (29) now serves as a general expression satisfying the momentum conservation and applicable for $m_{q}\neq m_{\bar{q}}$.
@@ IV.3 Numerical results
In Fig. 2, we plot $\rho^{\rm o}_{00}$ with $y_{q}$ dependence at fixed $\phi$ and $q_{T}$. Generically, the $y_{q}$ dependence is relatively mild. Basically, $\rho^{\rm o}_{00}$ monotonically increases for $q_{x}=q_{y}$ and decreases for $q_{y}=0$. For large $q_{T}$, $\rho^{\rm o}_{00}$ could be larger than $1/3$ for $q_{y}=0$ as already indicated Fig. 1. We further show contour plots for $\delta\rho^{\rm o}_{00}$ with respect to transverse momenta in Fig. 3. In general, $\rho^{\rm o}_{00}<1/3$ in most of kinematic regions.
@@
Analogously, we may also numerically evaluate $\rho^{\rm i}_{00}$ for the in-plane spin alignment for $\phi$ mesons. Due to symmetry, the results for $\rho^{\rm i}_{00}$ with $q_{T}$ dependence at fixed $\phi$ and $y_{q}$ are the same as those shown in Fig. 1 by exchanging the results for $\phi=0$ and $\phi=\pi/2$. In Fig. 4, we further present $\delta\rho^{\rm i}_{00}$ with $\phi$ dependence and $\rho^{\rm i}_{00}$ with $y_{q}$ dependence, respectively. In particular, for the large-momentum limit, we simply need to exchange $\hat{v}_{x}$ and $\hat{v}_{y}$ in Eq. (86), which yields $\delta\rho^{\rm o}_{00}\sim-(2-3\sin^{2}\phi)$, and the qualitatively features of $\delta\rho^{\rm i}_{00}$ with $\phi$ dependence at large $q_{T}$ are understood. Also, $\rho^{\rm i}_{00}<1/3$ for either $\phi=\pi$ or $\phi=0$ is expected. The $y_{q}$ dependence is generically mild. The contour plots in Fig. 5 correspond to those in Fig. 3 with $90$-degree rotations.
@@ V Spin alignment from color fields in QGP
$\rho^{\rm o}_{00}(q)=\frac{1-\big{[}\big{(}\tilde{C}_{B1}+\tilde{C}_{B2}\big{)}\chi^{yy}_{B}+\tilde{C}_{B1}(\chi^{xx}_{B}+\chi^{zz}_{B}\big{)}+\big{(}\tilde{C}_{E1}+\tilde{C}_{E2}\big{)}\chi^{yy}_{E}+\tilde{C}_{E1}(\chi^{xx}_{E}+\chi^{zz}_{E}\big{)}\big{]}}{3-\big{[}\big{(}3\tilde{C}_{B1}+\tilde{C}_{B2}\big{)}\chi_{B}+\big{(}3\tilde{C}_{E1}+\tilde{C}_{E2}\big{)}\chi_{E}\big{]}},$ (87)
@@ Acknowledgements.
The author would like to thank the hospitality of ECT* during the workshop “Spin and Quantum Features of QCD Plasma” and other participants, especially A. Tang, for helpful discussions. This work is supported by National Science and Technology Council (Taiwan) under Grants No. MOST 110-2112-M-001-070-MY3 and No. NSTC 113-2628-M-001-009-MY4 and by Academia Sinica under Project No. AS-CDA-114-M01.